# signature_mor.bib

### v1

@@ 3.3 Reduced order signature approximation
In detail, the reduced model (3.16) is obtained by removing the equation for $\hat{\mathcal{X}}_{t}^{2}$ in (3.12) and by setting $\hat{\mathcal{X}}_{t}^{2}=0$ in the dynamics of $\hat{\mathcal{X}}_{t}^{1}$. This is motivated by th minor relevance of $\hat{\mathcal{X}}_{t}^{2}$ leading to a reduced output $\widetilde{\mathcal{Y}}\approx\mathcal{Y}$. In particular, choosing $p=1$ and $z=\big{(}\begin{matrix}1&0&\dots&0\end{matrix}\big{)}^{\top}$, the reduced variable $\widetilde{\mathcal{X}}$ is a candidate for $\widetilde{\mathbb{X}}$ in (2.14).
@@ 4.1 Bergomi model
In the first example, we use a standard stochastic volatility model, namely the Bergomi model [bergomi2015stochastic], as our starting point. The Bergomi model is highly praised for its flexibility, and the ability to accurately fit equity markets, even with relatively few parameters. More concretely, the $n$-factor Bergomi model is given by
@@
with initial values $S_{0}$ for the stock price and an initial forward variance curve $\xi_{0}^{\cdot}$. Here, $Z,W^{1},\ldots,W^{n}$ are correlated standard Brownian motions, where $\rho$ denotes the correlation matrix of $(W^{1},\ldots,W^{n})$ and – following [guyon2022vix] – we denote the correlation between the Brownian motion $Z$ driving the stock price and the Brownian motion $W^{i}$ by $\rho_{Si}$.
@@
In the first numerical example, we take a two-factor Bergomi model, following the parameterization of [guyon2022vix, Section 3], which is a variant of the parameter Set I in [bergomi2015stochastic, p. 229] with constant initial forward variance curve. Specifically, we choose the parameters presented in Table 1. We here use the notation $\theta_{1}=w_{1}$, and the convention $w_{2}=1-\theta_{1}$.
@@
$3$ $2.63$ $0.42$ $0.69$ $0.7$ $-0.9$ $-0.9$ $1$ $0.04$
Table 1: Parameters for the Bergomi model used in our numerical example following [guyon2022vix, Section 3].
In our first numerical example, we take the Bergomi model as ground-truth, and proceed to
@@
    train a signature model in the sense of [cuchiero2022signature] with different truncation levels $m$, i.e., a model for the asset price of the form $S_{t}\approx S^{(\ell)}_{t}=\left\langle\ell\,,\mathbb{X}^{\leq m}_{0,t}\right\rangle$, see Section 2.6;
@@
    vectorize the truncated signature $\mathbb{X}^{\leq m}_{0,t}$ and formulate Itô-SDE (3.1) with respective initial state, so that $S^{(\ell)}_{t}=\mathcal{Y}_{t}=L\mathcal{X}_{t}$;
@@
    approximate the truncated signature by the method described in Section 3.3; obtain a reduced order signature approximation $S^{(\ell)}_{t}\approx\widetilde{\mathcal{Y}}_{t}=\widetilde{L}\widetilde{\mathcal{X}}_{t}$ from (3.16).
@@
Figure 1: Square root of first $50$ out of $n=1365$ eigenvalues of $PQ$ for signature model associated to (4.1).
@@
[…]
Figure 3: Implied volatilities of (4.1) for $T=1/12,1/2,1$ and strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ with $j=0,1,\dots,20$.
In addition to the $L^{2}$-error, the quality of the reduced signature system is tested in a finance context. First, we determine the fair price $\mathbb{E}\left[\max\{S^{(\ell)}_{T}-K_{T},0\}\right]$ of a European option, where the values of the strike price $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ ($j=0,1,\dots,20$) are chosen depending on the maturity $T$. The computed prices are then treated as prices from a Black-Scholes model and the associated volatilities are derived (while the interest rate is zero). These quantities are called implied volatilities (IV). The IV are depicted in Figure 3 for $T=1/12,1/2,1$.
@@
Figure 4: Relative error between volatilities of signature model associated to (4.1) and reduced systems with $\widetilde{n}=5,11,27$, strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ and $T=1/12,1/2,1$.
@@ 4.2 Rough Bergomi model
Our second example is the rough Bergomi model introduced in [bayer2016pricing], a work-horse model within the class of rough volatility models, see [bayer2023rough] for an exposition. Intuitively, the rough Bergomi model is a variant of the Bergomi model, where the exponential memory kernel in the variance process is replaced by a fractional kernel, in essence leading to the variance process being an exponential of a fractional Brownian motion. While seemingly innocuous, this change destroys the Markovian structure of the resulting model. More precisely, we consider the model
@@
The term “rough” in the rough Bergomi model reflects the empirical observation that the Hurst index $H$ should be chosen less than $\frac{1}{2}$ – leading to a power law explosion of the ATM implied volatility skew. In fact, empirical studies show that $H$ is often chosen very close to 00, e.g., $H=0.07$ reported in [bayer2016pricing] based on a calibration on SPX option prices as of February 4, 2010.
We again fit a signature model with underlying state process $X_{t}=(t,Z_{t},W_{t})$ (i.e. $d=3$) to a rough Bergomi model, taking model parameters reported in Table 2. The parameters are realistic, with the possible exception of our choice of $H=0.3$. While clearly within the range of Hurst parameters observed in the large scale study [bennedsen2022decoupling], it is comparatively large for rough volatility models. In our experience, fitting a signature model to a rough volatility model with very small $H$ is, however, hard, especially while keeping the truncation degree manageable. Therefore, we choose $H=0.3$ as a compromise of a clearly fractional, non-Markovian model which is still easily fitable by a signature model. Note that we again use a constant initial forward variance curve. In particular, we choose $m=7$ as the level of the truncated signature $\mathbb{X}^{\leq m}_{0,t}$ leading to an equation in (3.1) with state dimension $n=3280$. The reduction technique of Section 3 provides a reduced system (3.16) with state dimension $\widetilde{n}\ll n$ with the aim of having an accurate approximation of the quantity of interest, i.e., $\widetilde{\mathcal{Y}}_{t}\approx{\mathcal{Y}}_{t}$. The algebraic values $\sigma_{k}:=\sqrt{\operatorname{eig}_{k}(PQ)}$ in Figure 6 tell us about the significance of state variables and hence the right reduced dimension $\widetilde{n}$. We observe that $\sigma_{k}$ is numerically zero for $k>55$. Therefore, a reduced system of order $\widetilde{n}=55$ is an exact model. However, choosing $\widetilde{n}$ much smaller than this can go along with a little approximation error as well. We illustrate the $L^{2}$-performance for all case with a true error in Figure 6. We, e.g., observe that the $L^{2}$-error between $\widetilde{\mathcal{Y}}$ with $\widetilde{n}>10$ and ${\mathcal{Y}}$ (linear functional of the signature) is below $0.01$.
Figure 5: Square root of first $80$ out of $n=3280$ eigenvalues of $PQ$ for signature model associated to (4.2).
@@
Figure 7: Implied volatilities of (4.2) for $T=1/12,1/2,1$ and strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ with $j=0,1,\dots,20$.
@@
Figure 8: Relative error between volatilities of signature model associated to (4.2) and reduced systems with $\widetilde{n}=15,55$, strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ and $T=1/12,1/2,1$.
@@ Acknowledgments
MR is supported by the DFG via the individual grant “Low-order approximations for large-scale problems arising in the context of high-dimensional PDEs and spatially discretized SPDEs”– project number 499366908.

### v2

@@ 3.3 Reduced order signature approximation
In detail, the reduced model (3.17) is obtained by removing the equation for $\hat{\mathcal{X}}_{t}^{2}$ in (3.13) and by setting $\hat{\mathcal{X}}_{t}^{2}=0$ in the dynamics of $\hat{\mathcal{X}}_{t}^{1}$. This is motivated by th minor relevance of $\hat{\mathcal{X}}_{t}^{2}$ leading to a reduced output $\widetilde{\mathcal{Y}}\approx\mathcal{Y}$. In particular, choosing $p=1$ and $z=\big(\begin{matrix}1&0&\dots&0\end{matrix}\big)^{\top}$, the reduced variable $\widetilde{\mathcal{X}}$ is a candidate for $\widetilde{\mathbb{X}}$ in (A.12).
3.4 Algebraic error representation
It is our goal to approximate $\mathcal{Y}$ in (3.1) by the reduced output $\widetilde{\mathcal{Y}}$ defined in (3.17). It is essential to characterize the error of this approach without computing $\mathcal{Y}$, since this quantity is too computationally involved or not even available. Below, we derive an algebraic representation of the MOR error that can be computed and does not require any sampling of the underlying stochastic processes. First, we observe that
(3.18) $\int_{0}^{T}\mathbb{E}\|\mathcal{Y}_{t}-\widetilde{\mathcal{Y}}_{t}\|^{2}_{\mathbb{R}^{p}}\mathrm{d}t$ $=\mathbb{E}\int_{0}^{T}\big\|L\Phi(t,0)z-\widetilde{L}\widetilde{\Phi}(t,0)\widetilde{z}\big\|_{\mathbb{R}^{p}}\mathrm{d}t$
$=\mathbb{E}\int_{0}^{T}\big\|L^{e}\Phi^{e}(t,0)z^{e}\big\|_{\mathbb{R}^{p}}^{2}\mathrm{d}t$
$=\operatorname{tr}\big(L^{e}\int_{0}^{T}\mathbb{E}\big[\Phi^{e}(t,0)z^{e}{z^{e}}^{\top}\Phi^{e}(t,0)^{\top}\big]\mathrm{d}t\,{L^{e}}^{\top}\big)$
$=\operatorname{tr}\big(L^{e}\int_{0}^{T}Z^{e}_{t}\mathrm{d}t\,{L^{e}}^{\top}\big),$
where $Z^{e}_{t}=\mathbb{E}\left[\Phi^{e}(t,0)z^{e}{z^{e}}^{\top}{\Phi^{e}}^{\top}(t,0)\right]$, $\widetilde{\Phi}$ is the fundamental solution of the reduced model (3.17) and
[…]
We notice that $\Phi^{e}$ is the fundamental solution of a system like (3.1) with coefficients
$A^{e}=\left(\begin{matrix}A&0\\ 0&{\widetilde{A}}\end{matrix}\right),\quad N_{i}^{e}=\left(\begin{matrix}N_{i}&0\\ 0&{\widetilde{N}}_{i}\end{matrix}\right).$
Therefore, we know by Proposition 3.6 that $Z^{e}$ satisfies
(3.19) $\frac{\mathrm{d}}{\mathrm{d}t}Z^{e}_{t}=\mathcal{L}^{e}\left(Z^{e}_{t}\right),\quad Z^{e}_{0}=z^{e}{z^{e}}^{\top},\quad t\in[0,T],$
with $\mathcal{L}^{e}\left(Z^{e}\right):=A^{e}Z^{e}+Z^{e}{A^{e}}^{\top}+\sum_{i,j=2}^{d}N_{i}^{e}Z^{e}{N_{j}^{e}}^{\top}k_{ij}$. From (3.19), it can be seen that the left upper $n\times n$ block of $Z^{e}$ is the solution $Z$ of (3.7) with initial state $zz^{\top}$. Furthermore, the right lower $\widetilde{n}\times\widetilde{n}$ block $\widetilde{Z}$ and the right upper $n\times\widetilde{n}$ block $Z_{2}$ of $Z^{e}$ satisfy
(3.20) $\frac{\mathrm{d}}{\mathrm{d}t}\widetilde{Z}_{t}$ $=\widetilde{\mathcal{L}}\left(\widetilde{Z}_{t}\right),\quad\widetilde{Z}_{0}=\widetilde{z}\widetilde{z}^{\top},\quad t\in[0,T],$
(3.21) $\frac{\mathrm{d}}{\mathrm{d}t}Z_{2,t}$ $=\mathcal{L}_{2}\left(Z_{2,t}\right),\quad Z_{2,0}=z\widetilde{z}^{\top},\quad t\in[0,T],$
[…]
where $P$ is given by (3.3), $\widetilde{P}=\int_{0}^{T}\widetilde{Z}_{t}\mathrm{d}t$ and $P_{2}=\int_{0}^{T}Z_{2,t}\mathrm{d}t$. Identity (3.22) is now used to calculate the $L^{2}$-error of the MOR procedure. It is easily accessable, since $P$ is already available as a crucial object in the reduced system computation. Moreover, $\widetilde{P}$ and $P_{2}$ are integrals of the solutions of (3.20) and (3.21). Therefore, they can be computed analogue to $P$, see Section 3.2. On the other hand, $\widetilde{P}$ and $P_{2}$ are much easier to obtain as the are low-dimensional objects (depending on $\widetilde{n}$) in contrast to $P$.
@@ 4.1 Bergomi model
In the first example, we use a standard stochastic volatility model, namely the Bergomi model [14], as our starting point. The Bergomi model is highly praised for its flexibility, and the ability to accurately fit equity markets, even with relatively few parameters. More concretely, the $n$-factor Bergomi model is given by
@@
with initial values $S_{0}$ for the stock price and an initial forward variance curve $\xi_{0}^{\cdot}$. Here, $Z,W^{1},\ldots,W^{n}$ are correlated standard Brownian motions, where $\rho$ denotes the correlation matrix of $(W^{1},\ldots,W^{n})$ and – following [25] – we denote the correlation between the Brownian motion $Z$ driving the stock price and the Brownian motion $W^{i}$ by $\rho_{Si}$.
@@
In the first numerical example, we take a two-factor Bergomi model, following the parameterization of [25, Section 3], which is a variant of the parameter Set I in [14, p. 229] with constant initial forward variance curve. Specifically, we choose the parameters presented in Table 1. We here use the notation $\theta_{1}=w_{1}$, and the convention $w_{2}=1-\theta_{1}$.
@@
33 $2.63$ $0.42$ $0.69$ $0.7$ 0.9-0.9 0.9-0.9 11 $0.04$
Table 1: Parameters for the Bergomi model used in our numerical example following [25, Section 3].
In our first numerical example, we take the Bergomi model as ground-truth, and fit a signature model to it, which we then reduce following our MOR procedure. In more detail, we proceed according to the following steps.
@@
    First we train a signature model in the sense of [18] with different truncation levels $m$ minimizing the $L^{2}$-error between the terminal stock prices $\mathbb{E}\left[\left\lVert S_{T}-S^{\ell}_{T}\right\rVert^{2}\right]$ for $S^{(\ell)}_{t}=\left\langle\ell\,,\mathbb{X}^{\leq m}_{0,t}\right\rangle$, see Section A.6. Following the approach of [18], the resulting signature model is free of arbitrage.
@@
    We then vectorize the truncated signature $\mathbb{X}^{\leq m}_{0,t}$ and formulate the Itô-SDE (3.1) with respective initial state, so that $S^{(\ell)}_{t}=\mathcal{Y}_{t}=L\mathcal{X}_{t}$, for a suitable row-matrix $L$.
@@
    Finally, we approximate the truncated signature by the method described in Section 3.3 and obtain a reduced order signature approximation $S^{(\ell)}_{t}\approx\widetilde{\mathcal{Y}}_{t}=\widetilde{L}\widetilde{\mathcal{X}}_{t}$ from (3.17). As noted earlier, $\widetilde{\mathcal{X}}_{t}$ is simply the solution to a (low-dimensional) linear system of SDEs.
@@
Figure 1: Square root of first $50$ out of $n=1365$ eigenvalues of PQPQ for signature model associated with (4.1).
@@
Figure 2 shows $\sigma_{k}$ for $k\in\{1,\dots,50\}$. We observe a strong decay in $k$ and notice that $\sigma_{28}<10^{-8}$ which is below the machine precision. Consequently, (3.1) has a high reduction potential and allows for an exact approximation in case of choosing $\widetilde{n}=27$. The corresponding reduction errors $\sqrt{\mathbb{E}\int_{0}^{T}\left\|\mathcal{Y}_{t}-\widetilde{\mathcal{Y}}_{t}\right\|_{\mathbb{R}^{p}}^{2}dt}$ can be found in Figure 2. It is important to notice that this $L^{2}$-error is not calculated from a Monte-Carlo approach, but from the algebraic error representation in (3.22). We have a true approximation error for $\widetilde{n}<27$. If $\widetilde{n}\geq 27$, the error is numerically zero.
Figure 3: Implied volatilities of the signature-based approximation $S^{(\ell)}$ of (4.1) for $T=1/12,1/2,1$ and strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ with $j=0,1,\dots,20$.
In addition to the $L^{2}$-error, the quality of the reduced signature system is tested in a finance context. First, we determine the fair price $\mathbb{E}\left[\max\{S^{(\ell)}_{T}-K_{T},0\}\right]$ of a European option, where the values of the strike price $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ ($j=0,1,\dots,20$) are chosen depending on the maturity $T$. The computed prices are then treated as prices from a Black-Scholes model and the associated volatilities are derived (while the interest rate is zero). These quantities are called implied volatilities (IV). The IV of $S^{(\ell)}$ are depicted in Figure 3 for $T=1/12,1/2,1$.
@@
Figure 4: Relative error between volatilities of signature model associated with (4.1) and reduced systems with $\widetilde{n}=5,11,27$, strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ and $T=1/12,1/2,1$.
@@ 4.2 Rough Bergomi model
Our second example is the rough Bergomi model introduced in [3], a work-horse model within the class of rough volatility models, see [2] for an exposition. Intuitively, the rough Bergomi model is a variant of the Bergomi model, where the exponential memory kernel in the variance process is replaced by a fractional kernel, in essence leading to the variance process being an exponential of a fractional Brownian motion. While seemingly innocuous, this change destroys the Markovian structure of the resulting model. More precisely, we consider the model
@@
The term “rough” in the rough Bergomi model reflects the empirical observation that the Hurst index $H$ should be chosen less than $\frac{1}{2}$ – leading to a power law explosion of the ATM implied volatility skew. In fact, empirical studies show that $H$ is often chosen very close to 0, e.g., $H=0.07$ reported in [3] based on a calibration on SPX option prices as of February 4, 2010.
We again fit a signature model with underlying state process $X_{t}=(t,Z_{t},W_{t})$ (i.e. $d=3$) to a rough Bergomi model, taking model parameters reported in Table 2, using the same approach as in Section 4.1 for determining the signature model.
The parameters are realistic, with the possible exception of our choice of $H=0.3$. While clearly within the range of Hurst parameters observed in the large scale study [10], it is comparatively large for rough volatility models. In our experience, fitting a signature model to a rough volatility model with very small $H$ is, however, hard, especially while keeping the truncation degree manageable. Therefore, we choose $H=0.3$ as a compromise of a clearly fractional, non-Markovian model which is still easily fittable by a signature model. Note that we again use a constant initial forward variance curve. In particular, we choose $m=7$ as the level of the truncated signature $\mathbb{X}^{\leq m}_{0,t}$ leading to an equation in (3.1) with state dimension $n=3280$. The reduction technique of Section 3 provides a reduced system (3.17) with state dimension $\widetilde{n}\ll n$ with the aim of having an accurate approximation of the quantity of interest, i.e., $\widetilde{\mathcal{Y}}_{t}\approx{\mathcal{Y}}_{t}$. The algebraic values $\sigma_{k}:=\sqrt{\operatorname{eig}_{k}(PQ)}$ in Figure 6 tell us about the significance of state variables and hence the right reduced dimension $\widetilde{n}$. We observe that $\sigma_{k}$ is numerically zero for $k>55$. Therefore, a reduced system of order $\widetilde{n}=55$ is an exact model. However, choosing $\widetilde{n}$ much smaller than this can go along with a little approximation error as well. We illustrate the $L^{2}$-performance for all case with a true error in Figure 6. Once more, the $L^{2}$-error is computed from (3.22) and not from sampling the underlying processes. We, e.g., observe that the $L^{2}$-error between $\widetilde{\mathcal{Y}}$ with $\widetilde{n}>10$ and $\mathcal{Y}$} (linear functional of the signature) is below $0.01$.
Figure 5: Square root of first $80$ out of $n=3280$ eigenvalues of PQPQ for signature model associated with (4.2).
@@
Figure 7: Implied volatilities of the signature-based approximation $S^{(\ell)}$ of (4.2) for $T=1/12,1/2,1$ and strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ with $j=0,1,\dots,20$.
@@
Figure 8: Relative error between volatilities of signature model associated with (4.2) and reduced systems with $\widetilde{n}=15,55$, strike prices $K_{T}=(0.8+j\cdot 0.02)^{\sqrt{T}}$ and $T=1/12,1/2,1$.
@@ Acknowledgments
MR is supported by the DFG via the individual grant “Low-order approximations for large-scale problems arising in the context of high-dimensional PDEs and spatially discretized SPDEs” – project number 499366908 and via the SFB 1294, project A07 “ Data-based model order reduction for stochastic dynamics” – project number 318763901.
Appendix A Introduction to signatures
For convenience, this section offers a self-contained, conceptual summary of path signatures to support the main results. For more details on the underlying concepts of rough paths, we refer to [22] for a modern introduction as well as [23] for a classical monograph. More information on path signature methods specifically can be found in [7].
A.1 Signatures of smooth paths
(A.1) $\mathbbm{x}^{I}_{s,t}\coloneqq\int_{s<t_{1}<\cdots<t_{n}<t}\dot{x}^{i_{1}}_{t_{1}}\mathrm{d}t_{1}\cdots\dot{x}^{i_{n}}_{t_{n}}\mathrm{d}t_{n}\eqqcolon\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x^{i_{1}}_{t_{1}}\cdots\mathrm{d}x^{i_{n}}_{t_{n}}\in\mathbb{R},$
where $0\leq s\leq t\leq T$. Here, $x^{i}_{t}$ denotes the $i$th coordinate of the path $x$ evaluated at time $t$ for a fixed basis $e_{1},\ldots,e_{d}$ of $V$. We prefer to use the notation $\mathrm{d}x_{t}$” rather than $\dot{x}_{t}\mathrm{d}t$” because the former corresponds to the standard notation in the non-smooth case to be considered later – think of paths of Brownian motion.
(A.2) $\mathbbm{x}^{<\infty}_{s,t}\coloneqq 1+\sum_{n=1}^{\infty}\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x_{t_{1}}\otimes\cdots\otimes\mathrm{d}x_{t_{n}}\in T((V)),$
with the extended tensor algebra $T((V))$ being defined as
(A.3) $T((V))\coloneqq\prod_{n=0}^{\infty}V^{\otimes n}.$
Here, the initial term “11” is considered as the (sole) entry in $V^{\otimes 0}\simeq\mathbb{R}$. Note that elements of the extended tensor algebra $T((V))$ have infinitely many terms, and are, hence, comparable to formal power series – in $d$ non-commuting variables $e_{1},\ldots,e_{d}$. For numerical purposes, signatures need to be truncated at a finite degree. Hence, we also define
(A.4) $\mathbbm{x}^{\leq m}_{s,t}\coloneqq 1+\sum_{n=1}^{m}\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x_{t_{1}}\otimes\cdots\otimes\mathrm{d}x_{t_{n}}\in T^{m}(V)\coloneqq\bigoplus_{n=0}^{m}V^{\otimes n}.$
Both $T((V))$ and $T^{m}(V)$ are algebras under the tensor product $\otimes$ – the product of formal power series in non-commuting variables. We refer to [4] for more information. Using this product, we see that the signature (formally) solves a controlled ODE, namely $\dot{\mathbbm{x}}_{0,t}^{<\infty}=\mathbbm{x}_{0,t}^{<\infty}\otimes\dot{x}_{t}$ or (in a notation easier to adapt for non-smooth paths)
$\mathrm{d}\mathbbm{x}_{0,t}^{<\infty}=\mathbbm{x}_{0,t}^{<\infty}\otimes\mathrm{d}x_{t}=\sum_{i=1}^{d}\mathbbm{x}_{0,t}^{<\infty}\otimes e_{i}\mathrm{d}x^{i}_{t},\quad\mathbbm{x}_{0,0}^{<\infty}=1\in T((V)),$
where $e_{i}$, $i=1,\ldots d$, denote the standard basis vectors of $V$ as lifted to elements of $T((V))$, see also (2.1). The truncated signature satisfies the same ODE, but interpreted on the truncated tensor algebra $T^{m}(V)$, i.e., with a nilpotent tensor product, which will – abusing notation – still be denoted $\otimes$. That is, the truncated signature satisfies
$\mathrm{d}\mathbbm{x}_{0,t}^{\leq m}=\mathbbm{x}_{0,t}^{\leq m}\otimes\mathrm{d}x_{t}=\sum_{i=1}^{d}\mathbbm{x}_{0,t}^{\leq m}\otimes e_{i}\mathrm{d}x^{i}_{t},\quad\mathbbm{x}_{0,0}^{\leq m}=1\in T^{m}(V).$
Remark A.1.
We do not endow $T((V))$ with a topology, and, hence, the ODE (2.1) is only defined in a formal way. Several topologies for $T((V))$ have been considered in the rough path literature, including natural Hilbert or Banach sub-spaces of $T((V))$, see [17] for more details. In this paper, we will only ever use the full signature defined on $T((V))$ for motivation, the mathematical analysis will take place on $T^{m}(V)$ for finite truncation $m$. As a finite dimensional vector space, $T^{m}(V)$ will be equipped with the standard Euclidean metric.
The signature is invariant under re-parameterization of the path: I.e., if we consider a (smooth, increasing) map, say, $\gamma:[R,S]\to[0,T]$, and a path $\bar{x}:[R,S]\to V$, $u\mapsto x_{\gamma(u)}$, then $\mathbbm{x}_{\gamma(s),\gamma(t)}^{<\infty}=\bar{\mathbbm{x}}_{s,t}^{<\infty}$, $R\leq s\leq t\leq S$. Up to re-parameterization and so-called tree-like excursions, the signature $\mathbbm{x}^{<\infty}_{0,T}$ uniquely determines the path $x$, see [26]. Invariance under parameterization as well as the possibility of tree-like excursions can be avoided by adding the component $t$ to the path. We will often prefer to work with time-extended paths – i.e., paths with a component $x^{1}_{t}\equiv t$.
(A.5) $\left\langle\ell\,,\mathfrak{a}\right\rangle\coloneqq\sum_{j=1}^{k}\gamma_{j}\left\langle\mathbf{w}_{j}\,,\mathfrak{a}\right\rangle\coloneqq\sum_{j=1}^{k}\gamma_{j}\alpha_{(i^{j}_{1},\ldots,i^{j}_{K_{j}})}\in\mathbb{R}.$
A.2 Signatures for Hölder continuous paths
$\sup_{s<t}\frac{\|x_{t}-x_{s}\|}{|t-s|^{\alpha}}$
is finite. Then, we write $x\in C^{\alpha}$. We briefly sketch the concepts of rough paths and (truncated) signatures for $x\in C^{\alpha}$ – see [22] for more details. Formally, a rough path $\mathbf{x}$ is a two-parameter function from the simplex $\set{0\leq s\leq t\leq T}$ taking values in the truncated tensor algebra $T^{m}(V)$.
Remark A.2.
$\mathrm{d}y_{t}=U(y_{t})\mathrm{d}x_{t}.$
However, a formal Taylor expansion shows that higher order Euler approximations of the differential equation in terms of the iterated integral of $x$ of order up to $m=\left\lfloor 1/\alpha\right\rfloor$ would converge. Of course, the catch is that – once again – $C^{\alpha}$ paths are not regular enough for the iterated integrals to make sense in a classical way when $\alpha\leq 1/2$. On the flip side, if we enhance the path increments $x_{s,t}$ with higher order terms $\mathbbm{x}^{I}_{s,t}$ “behaving like iterated integrals” (see [22] for details), then we can solve the corresponding controlled differential equations for smooth enough vector fields in a pathwise, deterministic way.
$\varrho_{\alpha}(\mathbf{x},\widetilde{\mathbf{x}}):=\sum_{n=1}^{m}\sup_{0\leq s<t\leq T}\frac{\|\mathbbm{x}^{(n)}_{s,t}-\tilde{\mathbbm{x}}^{(n)}_{s,t}\|}{|t-s|^{n\alpha}},$
where $\mathbf{x}_{s,t}\coloneqq 1+\sum_{n=1}^{m}\mathbbm{x}^{(n)}_{s,t}\in T^{m}(V)$ and $\tilde{\mathbf{x}}$ is defined accordingly – with $\mathbbm{x}^{(1)}_{s,t}=x_{s,t}\in V$ and $\mathbbm{x}^{(n)}_{s,t}\in V^{\otimes n}$, $n=2,\ldots,m$, see [22, Def. 2.3] for the case $m=2$ and [23, Def. 8.6] for the general case. It is at the heart of rough path theory, and allows the extension of the theory of signatures – including universal approximation – from the bounded variation case to the semimartingale case by providing a path-wise integration theory as well as path-wise well–posedness of stochastic differential equations, we refer to the two references above for details. Now, we can define the desired concepts.
Definition A.3.
$\mathbf{x}_{s,t}=1+\sum_{n=1}^{N}\mathbbm{x}^{(n)}_{s,t}\in T^{N}(V)$